I need to find 2 equations to the following story problem so I can do the substitution method to solve. How many liters of a 30% alcohol solution must be mixed with 70 liters of a 90% solution to get a 50% solution? Can you show me the steps to figure this out. Thanks.
Answer by sofiyac(983)

lets say you need x liters of 30% solution, then y will be the liters of the 50% solution
x(30)+70*30=y(50)
x+70=y
so the two equations are
30x+2100=50y
x+70=y
solve using substitution